# RoomPainter: View-Integrated Diffusion for Consistent Indoor Scene Texturing

### v1

@@ Abstract
Indoor scene texture synthesis has garnered significant interest due to its important potential applications in virtual reality, digital media, and creative arts. Existing diffusion model-based researches either rely on per-view inpainting techniques, which are plagued by severe cross-view inconsistencies and conspicuous seams, or they resort to optimization-based approaches that entail substantial computational overhead. In this work, we present RoomPainter, a framework that seamlessly integrates efficiency and consistency to achieve high-fidelity texturing of indoor scenes. The core of RoomPainter features a zero-shot technique that effectively adapts a 2D diffusion model for 3D-consistent texture synthesis, along with a two-stage generation strategy that ensures both global and local consistency. Specifically, we introduce Attention-Guided Multi-View Integrated Sampling (MVIS) combined with a neighbor-integrated attention mechanism for zero-shot texture map generation. Using the MVIS, we firstly generate texture map for the entire room to ensure global consistency, then adopt its variant, namely Attention-Guided Multi-View Integrated Repaint sampling (MVRS) to repaint individual instances within the room, thereby further enhancing local consistency. Experiments demonstrate that RoomPainter achieves superior performance for indoor scene texture synthesis in visual quality, global consistency, and generation efficiency.
22footnotetext: Corresponding author.
@@ 1 Introduction
High-quality 3D content synthesis is in great demand, owing to its diverse applications across the entertainment industry, robotic simulation, and mixed-reality environments. With the advancement of text-to-image diffusion models [33, 34, 30, 28], large-scale indoor scene texture synthesis has made significant progress. Although existing indoor texture synthesis methods produce high-quality results, they face several significant challenges. Specifically, these methods can be categorized into two groups: inpainting-based methods [32, 5] and optimization-based methods [24, 6, 47]. On one hand, the inpainting-based methods utilize depth-to-image diffusion models [33, 51] to progressively generate textures for neighboring views, with the guidance of reference geometry and previously generated views. However, due to their per-view inpainting strategy, these approaches often lead to inconsistencies in texture style and semantics across non-adjacent views, and do not effectively address the loop-closure problem. On the other hand, optimization-based methods mitigate these inconsistencies by employing score distillation loss [29] or its variations [43] in global texture maps. However, the optimization process incurs significant time costs for indoor texture synthesis and faces training instability, which may introduce significant artifacts.
To address these challenges, we propose an efficient and high-fidelity indoor scene texturing framework, dubbed RoomPainter. The core of RoomPainter lies in a zero-shot technique that effectively adapts a 2D diffusion model for 3D-consistent texture synthesis, along with a two-stage generation strategy that ensures both global and local consistency. Specifically, we introduce Attention-Guided Multi-View Integrated Sampling (MVIS) combined with a related view-based attention mechanism for zero-shot texture map generation. Using the MVIS, we can generate a holistic texture map for the entire room, achieving better global consistency than per-view inpainting-based methods. Subsequently, to address the untextured areas raised by occlusions between objects while preserving the global consistency established in the first stage, we adopt a refine stage to decouple each instance in the room, then apply an Attention-Guided Multi-view Integrated Repaint Sampling (MVRS) technique at the instance scale to perform texture inpainting and refinement for each individual instance.
@@
  • We propose RoomPainter, an effective framework for indoor scene texture synthesis that excels in producing consistent, high-quality textures with remarkable time efficiency.
@@
  • We conduct extensive experiments and compare our method with strong baselines. The results demonstrate that our method not only achieves superior performance in visual quality for indoor texture synthesis but also excels in generation efficiency.
Figure 2: Illustration of Multi-view Integrated Sampling For $N$ viewpoints in the room, guided by the corresponding depth maps, we use a Diffusion model to generate the denoised observation $\mathcal{I}_{t}^{n}$ at timestep $t$. This observation is then projected into UV space to obtain the texture map for the respective viewpoint. The texture maps from multiple viewpoints are dynamically merged to produce the texture map for the current timestep, which subsequently guides the sampling process for the next timestep.
@@ 2 Related Works
With the rapid advancements in image [33, 28, 51, 26, 34] and video generation models [3, 44] in recent years, the field of 3D generation has experienced significant progress. Various methods now generate objects [29, 43, 8, 48, 41, 42] or scenes [23, 52, 49, 10] based on textual descriptions via leveraging SDS Loss and its variants [29, 43] to optimize implicit representation such as NeRF [25] and 3D Gaussian Splatting [19]. In the realm of scene generation, some approaches utilize text-guided inpainting models and depth prediction models, which incrementally construct scenes via a warping process [17, 11, 50, 27, 9]. Other methods focus on fine-tuning image generation models using 3D data, allowing them to generate multi-view consistent images or panorama for 3D tasks with the guidance of semantic images or depth maps. Moreover, some methods predict depth from a given image to initialize 3D Gaussian Splatting, followed by further optimization of the 3D Gaussian Splatting using video generation models.
@@
The goal of 3D object texturing methods is to create high-quality texture for the given 3D object that align with user-provided textual descriptions. Early methods, such as [32, 5, 39], utilized depth-aware Diffusion Models [33, 51, 26] to iteratively apply inpainting for each viewpoint. At each step, the texture is generated based on the texture from the previous viewpoint, gradually rotating the object to produce a complete texture. While these methods can generate textures in a relatively short time, they fail to address issues such as loop-closure and visible seams. Other works, such as [7, 24, 47], leverage differentiable rendering combined with SDS [29] to texture the given object through optimization techniques. Additionally, approaches like [4, 21, 13, 18] use UV maps as intermediaries and apply different fusion methods to integrate multiple of view independent UV maps into a consistent UV map. During the Diffusion model sampling process, this consistent UV map is used to adjust the Diffusion model’s input, ensuring that the resulting texture is more consistent, with fewer visible seams and improved multi-view consistency.
@@
In the task of 3D indoor scene texture generation, the objective is to create textures for a given indoor scene based on a user-provided textual description, which presents challenges such as object occlusion and multi-view consistency. SceneTex [6] defines an implicit texture field and optimizes it using VSD [43]. By employing an optimization-based approach to refine the texture field, SceneTex naturally resolves the issue of multi-view consistency in 3D scene textures. Furthermore, SceneTex incorporates a cross-attention mechanism to transfer textures from visible regions to occluded areas. However, this process can result in blurriness in the occluded regions, and the optimization procedure itself is computationally intensive. When methods like Text2Tex and TEXTure [5, 32], typically used for object texture generation, are applied to scene texture generation, they can only enforce style consistency across multiple objects based on textual control. However, these methods do not ensure global style consistency for the entire scene. Alternatively, if the entire room is treated as a single object and methods such as TEXTure or Text2Tex are employed with the camera positioned at the center of the room, generating the scene texture by rotating the camera, global style consistency can be achieved. However, this approach still faces challenges such as occlusion and visible seams between regions.
Other approaches, such as [45], use panoramic images to generate a rough texture for the entire room. The camera position is then changed to generate textures for occluded regions. Subsequently, an MLP is employed to predict the textures of unseen areas based on the color and texture coordinates from the UV map. However, this method requires the scene to have an initial texture and relies on texture style transfer, meaning it cannot generate textures from scratch for a completely untextured scene.
@@ 3 Methods
Given a 3D indoor scene mesh $\mathcal{M}$, our objective is to efficiently generate textures that not only align with the provided textual descriptions but also maintain coherent semantics. Since our framework employs pretrained T2I Diffusion model to generate textures for the given indoor scene mesh, we will start with an overview of the T2I Diffusion model and mesh rendering in Sec. 3.1. Following this, we introduce Multi-view Integrated Sampling (abbreviated as MVIS) in Sec. 3.2, which addresses the challenge of maintaining multi-view consistency while preserving high-frequency details in the diffusion model sampling process. Subsequently, In Sec. 3.3, we introduce a variant of MVIS, namely Multi-view Integrated Repaint Sampling (MVRS), to address the occlusion region and simultaneously refine existing textures. Finally, in Sec. 3.4, we introduce a Related View-based Attention mechanism used in the MVIS and MVRS processes to achieve more consistent image generation.
@@ 3.1 Preliminaries
Diffusion Model Sampling.  Diffusion models [16, 38] are latent variable models that comprises a forward process $q(x_{t}|x_{0})$ and a learned sampling process $p_{\theta}(x_{t-1}|x_{t})$. Given an initial data point $x_{0}$, its noisy counterpart at time t is denoted as $x_{t}$, and the forward process $q$ can be expressed as:
@@
where $t\in[0,\infty)$ and the values of $\alpha_{t}$ are defined by a scheduler to controls the amount of noise added at time step $t$. Additionally, it has $\beta_{t}=1-\alpha_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. In the sampling process $p_{\theta}$, with the DDPM sampler [16], the diffusion model progressively denoises an initial noise image $x_{T}\sim\mathcal{N}(0,I)$ to clean image. Specifically, at any intermediate time step $t$, the sampling process $p_{\theta}$ can be formed as:
@@
where $\epsilon_{\theta}(x_{t},t)$ is the estimated noise. As $t$ decreases from $T$ to 00, the sampling process $p_{\theta}$ ultimately samples a clean data point $x_{0}$. Moreover, this process can be generalized for learning a marginal distribution using an additional input condition. That leads depth-aware text-to-image diffusion models [1, 51], where the output of the model $\epsilon_{\theta}(x_{t},y,d,t)$ is conditioned on a text prompt $y$ and a depth map $d$.
In this paper, we utilize Stable Diffusion [28] as our base model. It performs denoising in the latent space and employs an autoencoder $\mathcal{D(\mathcal{E(\cdot)})}$ for the conversion between image and latent representations. Therefore, $x_{0}$ generated by its diffusion process will be decoded to the image $\mathcal{D}(x_{0})$ through its decoder.
@@
3.2 Multi-view Integrated Sampling
Different from object-level multi-view diffusion models [37, 20, 36], diffusion models that capable of ensuring multi-view consistency in scene level image generation [40] are extremely limited. To address this challenge, we propose Multi-view Integrated Sampling (MVIS), grounded with a Related View-based Attention technique, to adapt a standard T2I model to generate view-consistent images in a training-free manner. Based on MVIS, we first implement a global texturing stage to generate consistent textures for the scene mesh.
@@ 3.2 Multi-view Integrated Sampling
As shown in Algorithm. 1, in the sampling process, we firstly sample $\{x_{T}^{n}\}_{n=1}^{N}\sim\{\mathcal{N}({0},{I})\}$ for the $N$ camera views. Subsequently, in each denoising step $t$, we decode all $N$ estimated $x_{0,t}^{n}$ into $\mathcal{I}_{t}^{n}$ and back project them into individual per-view texture map $\mathcal{T}_{t}^{n}$. We then merge these per-view texture maps to form a global texture map $\mathcal{T}_{t}$ using dynamic merge [21, 2], which can be formulated as:
@@ 3.3 Multi-view Integrated Repaint Sampling
Figure 3: Illustration of Multi-view Integrated Repaint Sampling(MVRS) Due to occlusion between instances, some areas remain untextured after the first stage of texture generation. To address this issue, we perform texture generation for each instance in the room, conditioned on the painted areas. For the $N$ different viewpoints of each instance, guided by the corresponding depth maps, at sampling step $t$ of MVRS, we combine the painted areas: $x_{MVIS}$ with the sampling results of MVIS at step $t+1$: $x_{0,t+1}$ using a mask $P$. to form $x_{t}$, which serves as the input for the sampling process of the timestep t. Noise that corresponding to the current timestep was added before mask combine. Upon completing the MVRS sampling process, the textures for all instances in the room are fully generated.
@@
Specifically, we decouple each instance in the room, resulting a total of $K$ instances $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$. We then design a variant of MVIS, namely multi-view integrated repaint sampling (MVRS), to perform texture inpainting and refinement for each individual instance. As shown in Fig. 3, the core of MVRS lies in incorporating diffusion-based inpainting [22] into the MVIS process. For each individual instance, given $N$ predefined camera views, we firstly render $N$ images $\mathcal{I}^{n}$ along with their inpainting masks $P^{n}$ that indicates untextured region of them. Subsequently, for any time step $t$ in the MVIS process, we encode $\mathcal{I}^{n}$ into the latent space and add noise on it to produce a noisy latent $\tilde{x}_{t}^{n}$, followed by sampling ${x}_{t}^{n}$ from MVIS, then combine them into a new latent using the inpainting mask $P^{n}$, finally passing it into the U-Net for the next denoising step. This process introduces information from the painted area into the MVIS process, effectively transferring the texture of the painted area to the regions with holes while ensuring consistency across multiple viewpoints. After iterating through all instances within the room, we obtain the final output scene texture $\mathcal{T}_{MVRS}$.

### v2

@@ Abstract
Indoor scene texture synthesis has garnered significant interest due to its important potential applications in virtual reality, digital media and creative arts. Existing diffusion-model-based researches either rely on per-view inpainting techniques, which are plagued by severe cross-view inconsistencies and conspicuous seams, or adopt optimization-based approaches that involve substantial computational overhead. In this work, we present RoomPainter, a framework that seamlessly integrates efficiency and consistency to achieve high-fidelity texturing of indoor scenes. The core of RoomPainter features a zero-shot technique that effectively adapts a 2D diffusion model for 3D-consistent texture synthesis, along with a two-stage generation strategy that ensures both global and local consistency. Specifically, we introduce Attention-Guided Multi-View Integrated Sampling (MVIS) combined with a neighbor-integrated attention mechanism for zero-shot texture map generation. Using the MVIS, we firstly generate texture map for the entire room to ensure global consistency, then adopt its variant, namely Attention-Guided Multi-View Integrated Repaint Sampling (MVRS) to repaint individual instances within the room, thereby further enhancing local consistency and addressing the occlusion problem. Experiments demonstrate that RoomPainter achieves superior performance for indoor scene texture synthesis in visual quality, global consistency and generation efficiency.
@@ 1 Introduction
High-quality 3D content synthesis is in great demand, owing to its diverse applications across the entertainment industry, robotic simulation and mixed-reality environments. The design of indoor scene typically requires professionals to spend dozens of hours to complete. Generating texture for indoor scene through automated methods can significantly reduce labor costs and save time. With the advancement of text-to-image diffusion models [33, 34, 30, 28], large-scale indoor scene texture synthesis has made significant progress. Although existing indoor texture synthesis methods produce high-quality results, they face several significant challenges. Specifically, these methods can be categorized into two groups: inpainting-based methods [32, 6] and optimization-based methods [24, 7, 49]. Inpainting-based methods utilize depth-to-image diffusion models [33, 56] to progressively generate texture for neighboring views, with the guidance of reference geometry and previously generated views. However, due to their per-view inpainting strategy, these approaches often result in inconsistencies in texture style across instances within the room and fail to effectively address the loop-closure problem. Optimization-based methods mitigate these inconsistencies by employing score distillation loss [29] or its variations [43] in global texture maps. However, the optimization process incurs significant time costs for indoor scene texture synthesis, due to the introduction of a large number of learnable parameters and also prone to issue of training instability. To address these challenges, we propose an efficient and high-fidelity indoor scene texturing framework, dubbed RoomPainter. The core of RoomPainter lies in a zero-shot technique that effectively adapts a 2D diffusion model for 3D-consistent texture synthesis, along with a two-stage generation strategy that ensures both global and local consistency. Specifically, we introduce Attention-Guided Multi-View Integrated Sampling (MVIS) combined with a related view-based attention mechanism for zero-shot texture map generation. Using the MVIS, we generate a holistic texture map for the entire room, achieving better global consistency than per-view inpainting-based methods. Subsequently, to paint the untextured areas raised by occlusions between objects while preserving the global consistency established in the first stage, we adopt a refine stage to repaint each instance in the room. In this stage, we apply an Attention-Guided Multi-View Integrated Repaint Sampling (MVRS) technique at the instance scale to perform texture inpainting and refinement for each individual instance.
@@
  • We propose RoomPainter, an effective framework for indoor scene texture synthesis that excels in producing consistent, high-quality texture with remarkable time efficiency.
@@
  • We conduct extensive experiments and compare our method with strong baselines. The results demonstrate that our method not only achieves superior performance in visual quality for indoor scene texture synthesis but also excels in generation efficiency.
Figure 2: Illustration of Multi-view Integrated Sampling For $N$ surrounding viewpoints in the room, centered around the room’s center point and guided by the corresponding depth maps, we use a diffusion model to generate the denoised observation $\mathcal{I}_{t}^{n}$ at timestep $t$. This observation is then projected into UV space to obtain texture maps that corresponding to respective viewpoints. The texture maps from multiple viewpoints are dynamically merged to produce the texture map for the current timestep, which subsequently guides the sampling process for the next timestep.
@@ 2 Related Works
With the rapid advancements in image [33, 28, 56, 26, 34] and video generation models [3, 45, 4, 44, 54, 50] in recent years, the field of 3D generation has experienced significant progress. Various methods now generate objects [29, 43, 53, 11, 52, 51, 41] or scenes [18, 55, 27, 57, 54, 10] based on textual descriptions via leveraging SDS Loss and its variants [29, 43] to optimize implicit representation such as NeRF [25] and 3D Gaussian Splatting [20]. In the realm of scene generation, some approaches utilize text-guided inpainting models and depth prediction models, which incrementally construct scenes via a warping process [18, 12, 55, 27, 9]. Other methods focus on fine-tuning image generation models using 3D data, allowing them to generate multi-view consistent images or panorama for 3D tasks with the guidance of semantic images or depth maps. Moreover, some methods predict depth from a given image to initialize 3D Gaussian Splatting, followed by further optimization of the 3D Gaussian Splatting using video generation models.
@@
The goal of 3D object texturing methods is to create high-quality texture for the given 3D object that align with user-provided textual descriptions. Early methods, such as [32, 6, 39, 42], utilized depth-aware Diffusion Models [33, 56, 26] to iteratively apply inpainting for each viewpoint. At each step, the texture is generated based on the texture from the previous viewpoint, gradually rotating the object to produce a complete texture. While these methods can generate texture in a relatively short time, they fail to address issues such as loop-closure and visible seams. Other works, such as [8, 24, 49], leverage differentiable rendering combined with SDS [29] to texture the given object through optimization techniques. Additionally, approaches like [5, 22, 14, 19] use UV maps as intermediaries and apply different fusion methods to integrate multiple of view independent UV maps into a consistent UV map. During the Diffusion model sampling process, this consistent UV map is used to adjust the Diffusion model’s input, ensuring that the resulting texture is more consistent, with fewer visible seams and improved multi-view consistency.
@@
In the task of 3D indoor scene texture generation, the objective is to create texture for a given indoor scene based on a user-provided textual description, which presents challenges such as object occlusion and multi-view consistency. SceneTex [7] defines an implicit texture field and optimizes it using VSD [43]. By employing an optimization-based approach to refine the texture field, SceneTex naturally resolves the issue of multi-view consistency in 3D scene texture. Furthermore, SceneTex incorporates a cross-attention mechanism to transfer texture from visible regions to occluded areas. However, this process can result in blurriness in the occluded regions, and the optimization procedure itself is computationally intensive. When methods like Text2Tex and TEXTure [6, 32], typically used for object texture generation, are applied to scene texture generation, they can only enforce style consistency across multiple objects based on textual control. However, these methods do not ensure global style consistency for the entire scene. Alternatively, if the entire room is treated as a single object and methods such as TEXTure or Text2Tex are employed with the camera positioned at the center of the room, generating the scene texture by rotating the camera, global style consistency can be achieved. However, this approach still faces challenges such as occlusion and visible seams between regions.
Other approaches, such as [46], use panoramic images to generate a rough texture for the entire room. The camera position is then changed to generate texture for occluded regions. Subsequently, an MLP is employed to predict the texture of unseen areas based on the color and texture coordinates from the UV map. However, this method requires the scene to have an initial texture and relies on texture style transfer, meaning it cannot generate texture from scratch for a completely untextured scene.
@@ 3 Methods
Given a 3D indoor scene mesh $\mathcal{M}$, our objective is to efficiently generate texture that not only align with the provided textual descriptions but also maintain global consistency. Since our framework employs pretrained text-to-image diffusion model to generate texture for the given indoor scene mesh, we will start with an overview of the text-to-image Diffusion model and mesh rendering in Sec. 3.1. Following this, we introduce Multi-View Integrated Sampling (abbreviated as MVIS) in Sec. 3.2, which addresses the challenge of maintaining multi-view consistency while preserving high-frequency details in the diffusion model sampling process. Subsequently, In Sec. 3.3, we introduce a variant of MVIS, namely Multi-View Integrated Repaint Sampling (MVRS), to inpaint the occlusion region and simultaneously refine existing texture for each instance within the indoor scene. Finally, in Sec. 3.4, we introduce a Related View-based Attention mechanism used in the MVIS and MVRS processes, which enhances multi-view consistency during texture synthesis.
@@ 3.1 Preliminaries
Diffusion Model Sampling. Diffusion models [17, 38] are latent variable models that comprises a forward process $q(x_{t}|x_{0})$ and a learned sampling process $p_{\theta}(x_{t-1}|x_{t})$. Given an initial data point $x_{0}$, its noisy counterpart at time t is denoted as $x_{t}$, and the forward process $q$ is expressed as:
@@
where $t\in[0,\infty)$ and the values of $\alpha_{t}$ are defined by a scheduler to controls the amount of noise added to $x_{0}$ at time step $t$. Additionally, it has $\beta_{t}=1-\alpha_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. In the sampling process $p_{\theta}$, with the DDPM sampler [17], the diffusion model progressively denoises an initial noise image $x_{T}\sim\mathcal{N}(0,I)$ to clean image. Specifically, at any intermediate time step $t$, the sampling process $p_{\theta}$ can be formed as:
@@
where $\epsilon_{\theta}(x_{t},t)$ is the noise that estimated by the diffusion model. As $t$ decreases from $T$ to 00, the sampling process $p_{\theta}$ ultimately samples a clean data point $x_{0}$. Moreover, this process can be generalized for learning a marginal distribution using an additional input condition. That leads depth-aware text-to-image diffusion models [1, 56], where the output of the model $\epsilon_{\theta}(x_{t},y,d,t)$ is conditioned on a text prompt $y$ and a depth map $d$.
In this paper, we utilize Stable Diffusion [28] as our base model. It performs denoising in the latent space and employs an autoencoder $\mathcal{D(\mathcal{E(\cdot)})}$ for the conversion between image and latent representations. Therefore, latent $x_{0}$ generated by its diffusion process will be decoded to the image $\mathcal{D}(x_{0})$ through its decoder.
@@
3.2 Mult-View Integrated Sampling
Different from object-level multi-view diffusion models [37, 21, 36], diffusion models that capable of ensuring multi-view consistency in scene level image generation [40] are extremely limited.
To address this challenge, we propose Multi-View Integrated Sampling (MVIS), grounded with a Related View-based Attention technique, to adapt a standard text-to-image model to generate view-consistent images in a training-free manner. Based on MVIS, we first implement a global texturing stage to generate consistent texture for the scene mesh.
@@ 3.2 Mult-View Integrated Sampling
As shown in Algorithm. 1, in the sampling process, we firstly sample $\{x_{T}^{n}\}_{n=1}^{N}\sim\{\mathcal{N}({0},{I})\}$ for $N$ camera views. Subsequently, at each denoising step $t$, we decode all $N$ estimated $x_{0,t}^{n}$ into $\mathcal{I}_{t}^{n}$ and back project them into independent per-view texture map $\mathcal{T}_{t}^{n}$. We then merge these per-view texture maps to form a global texture map $\mathcal{T}_{t}$ using dynamic merge [22, 2], which can be formulated as:
@@ 3.3 Multi-View Integrated Repaint Sampling
Figure 3: Illustration of Multi-View Integrated Repaint Sampling(MVRS) Due to occlusion between instances, certain areas remain untextured after the first stage of texture generation. To address this issue, we perform texture generation for each instance within the room based on the painted areas. For $N$ different viewpoints of each instance, guided by the corresponding depth maps, at sampling step $t$ of MVRS, we combine the painted areas($x_{MVIS}$) with the sampling results of MVIS at step $t+1$ ($x_{0,t+1}$) using a mask $P$ to form $x_{t}$, which serves as the input for the sampling process of timestep t. Noise corresponding to the current timestep was added before mask combine. Upon completing the MVRS process, the texture map for a specific instance in the room is fully generated.
@@
Specifically, we decouple each instance in the room, resulting a total of $K$ instances $\mathcal{M}_{1},\dots,\mathcal{M}_{K}$. We then design a variant of MVIS, namely multi-view integrated repaint sampling (MVRS), to perform texture inpainting and refinement for each individual instance. As shown in Fig. 3, the core of MVRS lies in incorporating diffusion-based inpainting [23] into the MVIS process. For each individual instance, given $N$ predefined camera views, we firstly render $N$ images $\mathcal{I}^{n}$ along with their inpainting masks $P^{n}$ that indicates untextured region of them. Subsequently, for any time step $t$ in the MVIS process, we encode $\mathcal{I}^{n}$ into the latent space and add noise on it to produce a noisy latent $\tilde{x}_{t}^{n}$, followed by sampling ${x}_{t}^{n}$ from MVIS, then combine them into a new latent using the inpainting mask $P^{n}$, where the painted area keep unchanged, finally passing it into the U-Net for the next denoising step. This process introduces information from the painted area into the MVIS process, effectively transferring the texture of the painted area to the regions with holes while ensuring consistency across multiple viewpoints. After iterating through all instances within the room, we obtain the final output scene texture $\mathcal{T}_{MVRS}$.